# Explainable Natural Language Processing for Corporate Sustainability Analysis

### v2

@@ 1 Introduction
In this work, we focus on sustainability analyses of entities that provide sustainability reports (also referred to as sustainability disclosures); in particular public companies and institutions. These entities are crucial components in the global development of sustainability [14], and we refer to their sustainability as corporate sustainability.
Given these entities’ substantial stake in global sustainability, the analysis of their sustainability, or corporate sustainability analysis, is critical [70]. Yet, this is extremely challenging due to the inherent complexity of corporate sustainability as a concept. To elaborate, the complexities within firms (e.g., geography, size, business activities) and outside the firms (e.g., their global supply chains) as well as the firms’ relationships with non-business stakeholders lead to the evolving frameworks and guidelines put forward to address corporate sustainability issues [11][24]. Further complicating the matter is the lack of globally mandated standards for sustainability reporting that guides companies’ efforts in determining the extent and nature of their sustainability disclosures [5].
For example, a firm may opt to use a convenient corporate sustainability reporting framework that magnifies only certain aspects of a firm’s practices [20], while omitting key sustainability dimensions [23]. Or, within a given regulatory disclosure framework that a firm shall abide to, it may provide sustainability disclosures that are difficult for external parties to digest, and whose integrity and transparency may be questionable [10].
@@
Indeed, today’s corporate sustainability analysis is carried out by human analysts who read sustainability disclosures provided by companies (i.e., the data dimension), and provide corporate sustainability evaluations based on a combination of sustainability frameworks (e.g., Environmental, Social, Governance (ESG) [26], Global Reporting Initiative (GRI) [51], Sustainability Accounting Standards Board (SASB) [30], Greenhouse Gas Protocol and Carbon Disclosure Project (CDP) [65], United Nations’ Sustainable Development Goals (SDG) [67], United Nation Global Compact (UNGC) [62]). These evaluations require an extensive amount of human hours [59], and yet are inevitably influenced by the inherent biases in the frameworks adopted as well as the analysts’ own subjectivity [12][41].
This second aspect, that we label the analyst dimension, is also inherently subjective, similarly to the data dimension.
@@ 2 The subjectivity challenge
As a result of the complexity inherent to corporate sustainability, subjectivity commonly underpins corporate sustainability analysis. Specific challenges arising from subjectivity can be distilled into the data and analyst dimensions. We expound on these dimensions to frame the discourse on (X)NLP’s suitability for corporate sustainability analysis.
@@ 2.1 The data dimension
Additional data may come from government-mandated sustainability disclosures, where companies are required to release specific aspects of their sustainability performance [80], such as their environmental impact [75], or ESG (Environmental, Social, and Governance) practices [44]. Data originating from NGO sources and media coverage may also be relevant to assess a company’s sustainability efforts [86][4]. Yet these data are influenced by sustainability disclosures [6][33], underlining the importance of focusing our discussion on the sustainability disclosures themselves.
Due to the intricacies and multi-faceted dimensions of corporate sustainability, these data sources inherently contain subjectivity. They involve competing notions of corporate sustainability, expressed through a vast and diverse array of sustainability frameworks [71][57]. Moreover, even within a specific framework, deciding what information to convey necessitates a value judgement (e.g., financial vs. environmental materiality), further exacerbating the partiality involved [7]. We argue that this subjectivity at the data level manifests itself in four data grand challenges that analysts need to face, namely incompleteness, unreliability, ambiguity, and sophistication. These 4 data challenges are reported on the first block of (Fig. 1(A)), referred to as data dimension. We detail them in the following.
@@ 2.2 The analyst dimension
Limited human-hours. Analysts have a limited amount of time to produce corporate sustainability assessments. However, the process entails extracting meaningful information from corporate sustainability disclosures [36], which requires a significant amount of time i.e., human-hours [59]. In particular, the disclosures are extremely time consuming to read and understand due to the sophistication and ambiguity of the data within them. Moreover, given the potential incompleteness of disclosure information, analysts may analyse additional sources to ascertain a firm’s sustainability, further increasing the human-hours required [68].
Potential Bias. Analysts’ interpretations of sustainability data are subject to bias. This is not least due to the complex and wide ranging factors to be considered, such as firm characteristics, evolving sustainability guidelines, and economic implications, amongst others, often leading to leading to varying interpretations amongst analysts [12][41]. To further compound the potential for bias, data ambiguity and unreliability hinders clear interpretation of sustainability information by obfuscating negative content [31].
@@ 2.3 NLP to the rescue
Subjectivity pervades through the data and analyst dimensions, making corporate sustainability analyses an onerous task. To alleviate this burden, we argue that AI technologies, through linguistic understanding enabled by natural language processing (NLP), can significantly aid analysts. NLP can automatically process hundreds of sustainability disclosures [79], summarising key details and extracting insights for analysts. This reduces the human hours required to analyse disclosures and partially addresses potential bias by providing more consistent and replicable insights [2][17]. We concretise NLP’s usefulness in this domain by detailing five NLP tasks for corporate sustainability analysis (T1-T5) that we deem critical, and that have been partially explored in the literature.
@@
(T2) Sentiment Classification. Sentiment insights can also be derived from sustainability disclosures, providing analysts a valuable perspective on corporate sustainability. Polarity classification determines the emotional polarity of text from sustainability disclosures. For instance, within SEC 10k reports, sustainability related sentences can be labelled as positive, negative or neutral [63]. Alternatively, strategic classification uncovers a text’s prospective impact. Examples include the risk and opportunity impact analysis of ESG-related texts [82].
(T3) Rating. A firm’s sustainability performance can be automatically scored to expedite an analyst’s evaluation process. Aggregated rating methods include computing the frequency of sentiment labels for ESG-related headlines [34], and averaging the E, S, and G topic classification probabilities of documents [78]. Direct methods explicitly score companies without amalgamating already labelled data. For instance, the prediction of ESG risk ratings from sustainability topics [46]. Outside of holistic sustainability frameworks (i.e. ESG), direct methods also include estimating carbon emissions from bank transactions.
(T4) Dialogue. Processing sustainability disclosures into interactive dialogue format allows an analyst a user-friendly way for querying information. For instance, a chatbot can be queried to summarise or answer questions about a disclosure [59].
@@ 3.2 Semantic methods
As they attain a high level of semantic understanding, BERT models are suitable for tasks within the corporate sustainability analysis field that require complex language understanding, without being specifically structured for generative tasks. Therefore, BERT is appropriate for complex language understanding tasks like (T1) topic extraction, (T2) sentiment classification and (T3) rating, but not generative tasks such as (T4) dialogue. In line with this, BERT-like models have been used for ESG label classification [47], producing company environmental scores [55], and to derive topics and their associated keywords for ESG-related risk factors [46] using BERT-generated text embeddings.
@@ 3.3 Syntactic methods
Syntactic methods include part-of-speech (POS) tagging [15], dependency parsing [48], and constituency parsing [87], amongst others. These methods can augment efforts to accomplish several complex NLP tasks for corporate sustainability analysis, although they have only seen limited application so far. For instance, when combined with other methods like semantic analysis, text can be analysed more granularly. In other domains such as finance and economics, aspect-based sentiment analysis (ABSA) has already been achieved through POS tagging and semantic rules [54, 19]. Such integrated approaches can more intricately tackle task (T2) sentiment classification, by revealing sentiment insights toward aspects. To illustrate with an example, within the sentence “Company X has good emissions performance despite having poor employee wellness”, ABSA allows for the aspect ‘emissions performance’ to be labelled as positive and the aspect ‘employee wellness’ to be labelled as negative. Distinguishing between the sentiment of the two aspects enables a finer analysis of sustainability text. Additionally, syntactic analysis can also be used for identifying companies and their relationships with respect to sustainability activities [29]. Such methodologies can be enabled by explicit syntactic attributes (i.e., POS, dependency parsing) in addition to semantic patterns. Topics can be derived from the extracted links between companies and their sustainability practices to solve task (T1) topic extraction. Additionally, these links can be further processed to evaluate corporate sustainability performance, in the spirit of task (T3) rating. Other applications of syntactic analysis include evaluating the grammatical clauses of disclosure sentences [18], in line with (T5) linguistic patterns, or the parsing of sustainability-related concepts, in a similar vein to the extraction of financial concepts [28].
Figure 2: Alongside outputting a decision for sustainability sentiment classification, XAI can provide an interpretation and explanation
@@ 4 XNLP enhances subjectivity mitigation
Faithfulness. The accuracy of an AI model’s interpretability and explainability with respect to its true workings [50, 83, 85]. For example, the extent to which a model’s feature salience scores and provided explanations are representative of its actual workings.
@@ 4.1 Achieving trustworthy insights
For example, figure 2 highlights how interpretability, explainability and faithfulness enhances trust in an NLP model’s classification for sustainability text. Interpretability allows an analyst to verify that appropriate and relevant text features (i.e. ‘rearing animals’ instead of ‘portfolios’), are exploited for classifying the sentence’s risk sentiment (similar to the concepts used in self-interpretable image classification – e.g., [84]). Explainability explicates the relevance of ‘rearing animals’ to sustainability risk, further justifying the model’s classification. It explains the link between ‘rearing animals’, deforestation, methane emissions, and strained water resources, resolving the ambiguity surrounding the term ‘rearing animals’. Faithfulness ensures that the insights from interpretability and explainability are meaningful in that they accurately reflect the model’s mechanism. Put together, interpretability, explainability and faithfulness prove the robustness of the model’s decision-making process, increasing an analyst’s trust in its classification output.
@@ 4.3 Generalisability across frameworks and sources
Sustainability disclosures stem from different sources and are presented through diverse frameworks. As such, they can have differing linguistic and textual features, posing a challenge for NLP models to generalise. To elaborate, while firms can focus corporate sustainability disclosures on positive efforts [16], government-linked sustainability disclosures may mandate firms to disclose sustainability related risks (such as U.S. Securities and Exchange Commission 10k reports) [45]. A sentiment classification model trained on corporate sustainability disclosures may not adapt well to government-linked sustainability disclosures, as the latter more frequently carries negative sentiment. On top of heterogeneous sustainability data, the generalisability problem is further compounded by the lack of available datasets within the field [47]. This reduces the diversity of data for training robust models. In light of limited data availability, interpretability methods provides an alternative means for mitigating generalisability issues.

### v3

@@ 1 Introduction
In this work, we focus on sustainability analyses of entities that provide sustainability reports (also referred to as sustainability disclosures); in particular public companies and institutions. These entities are crucial components in the sustainable development [14], and we refer to their sustainability as corporate sustainability.
Given that these entities have a substantial stake in global sustainability, the analysis of their sustainability, or corporate sustainability analysis, is critical [73]. Yet, this is extremely challenging due to the inherent complexity of corporate sustainability as a concept. To elaborate, the complexities within firms (e.g., geography, size, business activities) and outside the firms (e.g., their global supply chains), as well as the firms’ relationships with non-business stakeholders, lead to the evolving frameworks and guidelines put forward to address corporate sustainability issues [11, 24]. Further complicating the matter is the lack of globally mandated standards for sustainability reporting [5].
For example, a firm may use a convenient corporate sustainability reporting framework that magnifies only certain aspects of a firm’s practices [20], while omitting key sustainability dimensions [23]. Or, within a given regulatory disclosure framework that a firm shall abide to, it may provide sustainability disclosures that are difficult for external parties to digest, and whose integrity and transparency may be questionable [10].
@@
Indeed, today’s corporate sustainability analysis is carried out by human analysts who read sustainability disclosures provided by companies (i.e., the data dimension), and provide corporate sustainability evaluations based on a combination of sustainability frameworks (e.g., Environmental, Social, Governance (ESG) [26], Global Reporting Initiative (GRI) [54], Sustainability Accounting Standards Board (SASB) [30], Greenhouse Gas Protocol and Carbon Disclosure Project (CDP) [68], United Nations’ Sustainable Development Goals (SDG) [70], United Nation Global Compact (UNGC) [65]). These evaluations require an extensive amount of human hours [62], and yet are inevitably influenced by the inherent biases in the frameworks adopted as well as the analysts’ own subjectivity [12, 42].
This second aspect, that we label the analyst dimension, is also inherently subjective, similar to the data dimension.
@@ 2 The subjectivity challenge
As a result of the complexity inherent to corporate sustainability, subjectivity commonly underpins corporate sustainability analysis. Specific challenges arising from subjectivity can be distilled into the data and analyst dimensions. We expound on these dimensions to frame (X)NLP’s suitability for corporate sustainability analysis.
@@ 2.1 The data dimension
Additional data may come from government-mandated sustainability disclosures, where companies are required to release specific aspects of their sustainability performance [84], such as their environmental impact [78], or ESG (Environmental, Social, and Governance) practices [47]. Data originating from NGO sources and media coverage may also be relevant to assess a company’s sustainability efforts [91, 4]. Yet these data are influenced by sustainability disclosures [6, 33], underlining the importance of focusing our discussion on the sustainability disclosures themselves.
Due to the intricacies and multi-faceted dimensions of corporate sustainability, these data sources inherently contain subjectivity. They involve competing notions of corporate sustainability, expressed through a vast and diverse array of sustainability frameworks [74, 60]. Moreover, even within a specific framework, deciding what information to convey necessitates a value judgement (e.g., financial vs. environmental materiality), further exacerbating the partiality involved [7]. We argue that this subjectivity at the data level manifests itself in four data grand challenges that analysts need to face, namely incompleteness, unreliability, ambiguity, and sophistication. These 4 data challenges are reported on the first block of (Fig. 1(A)), referred to as data dimension. We detail them in the following.
@@ 2.2 The analyst dimension
Limited human-hours.
Analysts have a limited amount of time to produce corporate sustainability assessments. However, the process entails extracting meaningful information from corporate sustainability disclosures [37], which requires a significant amount of time i.e., human-hours [62]. In particular, the disclosures are extremely time consuming to read and understand due to the sophistication and ambiguity of the data within them. Moreover, given the potential incompleteness of disclosure information, analysts may analyse additional sources to ascertain a firm’s sustainability, further increasing the human-hours required [71].
Potential Bias.
Analysts’ interpretations of sustainability data are subject to bias. This is not least due to the complex and wide ranging factors to be considered, such as firm characteristics, evolving sustainability guidelines, and economic implications, amongst others, often leading to leading to varying interpretations amongst analysts [12, 42]. To further compound the potential for bias, data ambiguity and unreliability hinders clear interpretation of sustainability information by obfuscating negative content [31].
@@ 2.3 NLP to the rescue
Subjectivity pervades through the data and analyst dimensions, making corporate sustainability analyses an onerous task. To alleviate this burden, we argue that AI technologies, through linguistic understanding enabled by natural language processing (NLP), can significantly aid analysts. NLP can automatically process hundreds of sustainability disclosures [83], summarising key details and extracting insights for analysts. This reduces the human hours required to analyse disclosures and partially addresses potential bias by providing more consistent and replicable insights [2, 17]. We concretise NLP’s usefulness in this domain by detailing five NLP tasks for corporate sustainability analysis (T1-T5) that we deem critical, and that have been partially explored in the literature.
@@
(T2) Impact Classification. Impact insights can also be derived from sustainability disclosures, providing analysts a valuable perspective on corporate sustainability. Polarity classification determines the impact polarity of text from sustainability disclosures. For instance, within SEC 10k reports, sustainability related sentences can be labelled as positive, negative or neutral in relation to their impact toward sustainability [66]. Alternatively, strategic classification uncovers a text’s prospective impact. Examples include the risk and opportunity impact analysis of ESG-related texts [87].
(T3) Rating. A firm’s sustainability performance can be automatically scored to expedite an analyst’s evaluation process. Aggregated rating methods include computing the frequency of sentiment labels for ESG-related headlines [34], and averaging the E, S, and G topic classification probabilities of documents [82]. Direct methods explicitly score companies without amalgamating already labelled data. For instance, the prediction of ESG risk ratings from sustainability topics [49], as well as regressing lexical features to derive ESG risk ratings [44]. Outside of holistic sustainability frameworks (i.e. ESG), direct methods also include estimating carbon emissions from bank transactions [38], and constructing sustainability indices from the lexical features of disclosures [85].
(T4) Dialogue. Processing sustainability disclosures into interactive dialogue format allows an analyst a user-friendly way for querying information. For instance, a chatbot integrated with Retrieval-augmented Generation (RAG) can be queried to summarise or answer analysts’ questions about disclosures, allowing analysts to extract information without reading the copious amounts of text typically contained within them [62, 36]. By emphasising critical information, dialogue systems mitigate the obfuscation of key information within disclosures, potentially enhancing the accuracy of sustainability analysis [79].
@@ 3.2 Semantic methods
As they attain a high level of semantic understanding, BERT models are suitable for tasks within the corporate sustainability analysis field that require complex language understanding, without being specifically structured for generative tasks. Therefore, BERT is appropriate for complex language understanding tasks like (T1) topic extraction, (T2) impact classification and (T3) rating, but not generative tasks such as (T4) dialogue. In line with this, BERT-like models have been used for ESG label classification [50], producing company environmental scores [58], and to derive topics and their associated keywords for ESG-related risk factors [49] using BERT-generated text embeddings.
@@ 3.3 Syntactic methods
Syntactic methods include part-of-speech (POS) tagging [15], dependency parsing [51], and constituency parsing [92], amongst others. These methods can augment efforts to accomplish several complex NLP tasks for corporate sustainability analysis, although they have only seen limited application so far. For instance, when combined with other methods like semantic analysis, text can be analysed more granularly. In other domains such as finance and economics, aspect-based sentiment analysis (ABSA) has already been achieved through POS tagging and semantic rules [57, 19]. Such integrated approaches can more intricately tackle task (T2) impact classification, by revealing impact insights toward aspects. To illustrate with an example, within the sentence “Company X has good emissions performance despite having poor employee wellness”, ABSA allows for the aspect ‘emissions performance’ to be labelled as positive and the aspect ‘employee wellness’ to be labelled as negative. Distinguishing between the impact of the two aspects enables a finer analysis of sustainability text. Additionally, syntactic analysis can also be used for identifying companies and their relationships with respect to sustainability activities [29]. Such methodologies can be enabled by explicit syntactic attributes (i.e., POS, dependency parsing) in addition to semantic patterns. Topics can be derived from the extracted links between companies and their sustainability practices to solve task (T1) topic extraction. Additionally, these links can be further processed to evaluate corporate sustainability performance, in the spirit of task (T3) rating. Other applications of syntactic analysis include evaluating the grammatical clauses of disclosure sentences [18], in line with (T5) linguistic patterns, or the parsing of sustainability-related concepts, in a similar vein to the extraction of financial concepts [28].
Figure 2: Alongside outputting a decision for sustainability impact classification, XAI can provide an interpretation and explanation
@@ 4 XNLP enhances subjectivity mitigation
Faithfulness. The accuracy of an AI model’s interpretability and explainability with respect to its true workings [53, 88, 90]. For example, the extent to which a model’s feature salience scores and provided explanations are representative of its actual workings. Analysing faithfulness can be challenging given the lack of ground truth for a model’s explainability or interpretability, with several approaches involving dataset modification to verify the feature importance estimated by interpretability methods [43].
@@ 4.1 Achieving trustworthy insights
For example, figure 2 highlights how interpretability, explainability and faithfulness enhances trust in an NLP model’s classification for sustainability text. Interpretability allows an analyst to verify that appropriate and relevant text features (i.e. ‘rearing animals’ instead of ‘portfolios’), are exploited for classifying the sentence’s risk impact (similar to the concepts used in self-interpretable image classification – e.g., [89]). Explainability explicates the relevance of ‘rearing animals’ to sustainability risk, further justifying the model’s classification. It explains the link between ‘rearing animals’, deforestation, methane emissions, and strained water resources, resolving the ambiguity surrounding the term ‘rearing animals’. Faithfulness ensures that the insights from interpretability and explainability are meaningful in that they accurately reflect the model’s mechanism. Put together, interpretability, explainability and faithfulness prove the robustness of the model’s decision-making process, increasing an analyst’s trust in its classification output.
@@ 4.3 Generalisability across frameworks and sources
Sustainability disclosures stem from different sources and are presented through diverse frameworks. As such, they can have differing linguistic and textual features, posing a challenge for NLP models to generalise. To elaborate, while firms can focus corporate sustainability disclosures on positive efforts [16], government-linked sustainability disclosures may mandate firms to disclose sustainability related risks (such as U.S. Securities and Exchange Commission 10k reports) [48]. An impact classification model trained on corporate sustainability disclosures may not adapt well to government-linked sustainability disclosures, as the latter more frequently carries negative impact. On top of heterogeneous sustainability data, the generalisability problem is further compounded by the lack of available datasets within the field [50]. This reduces the diversity of data for training robust models. In light of limited data availability, interpretability methods provides an alternative means for mitigating generalisability issues.
@@ 4.4 Mitigating bias
5 Conclusion
XNLP algorithms are well equipped for sustainability analysis given that they encompass both linguistic understanding and XAI capabilities. To elaborate on this, while lexical, semantic and syntactic analysis of text automates and enhances sustainability analysis, XAI further mitigates the subjectivity challenge of sustainability assessments. However, NLP for sustainability analysis has not yet embraced these directions, leaving its vast potential untapped. This underscores an opportunity for researchers and developers to explore the strategies highlighted—achieving trustworthy insights, inferring patterns for greenwashing research, generalisability across frameworks and sources, mitigating bias—to enhance NLP for tasks (T1-T5). These efforts can not only improve sustainability analysis but potentially reshape how businesses, investors, and regulators address corporate sustainability reporting and attribution.
Limitations
Given that the development of NLP methods for sustainability analysis is still nascent, several papers that describe NLP methods for sustainability analysis entail emerging research that have not yet undergone extensive peer-review. Despite our attempts to filter out papers with evident shortcomings, it is important to acknowledge that some of the discussed papers may contain non-authoritative viewpoints.